OpenStudy (anonymous):

Silly seeming question. (I'm really bad at math) 6x/16x = 3/8x or 3/8?

OpenStudy (anonymous):

i would think 3/8... ?

OpenStudy (ahsome):

Not a silly question ;) Equation is: \[\frac{6x}{16x}\] Simplify by dividing both by \(3x\) \[\frac{3}{8}\] Thats the simplified version. Answer is \(\frac{3}{8}\)

OpenStudy (anonymous):

! I was right!! Great!! Thanks!

OpenStudy (ahsome):

No problems @CheyenneC :D